If equations of a projectile projected at angle θ with horizontal are following

x = 15t (m)

y = 15t – 5t2 (m), (where x and y are co-ordinates) then speed after 1 second of projection is
1
Expert's answer
2018-09-17T15:31:08-0400
The velocity components of a projectile
v_x (t)=15 (m/s)
v_y (t)=15-10t (m/s)
Therefore, speed after t=1 second of projection
v(1)=√(v_x^2 (1)+v_y^2 (1))=√(〖15〗^2+(15-10)^2)=15.8 m/s≅16 m/s